OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y < -3x + 3 y < x + 2

OpenStudy (texaschic101):

answer choices please

OpenStudy (anonymous):

oh sry i forgot

OpenStudy (anonymous):

(1, -5) (1, 5) (5, 1) (-1, 5)

OpenStudy (texaschic101):

just a matter of subbing in your points to see which ones are true. The solution set will make each line correct. y < -3x + 3 y < x + 2 (1,-5)...x = 1 and y = -5 -5 < -3(1) + 3 -5 < -3 + 3 -5 < 0 (true) -5 < 1 + 2 -5 < 3 (true) solution is (1,-5)

OpenStudy (anonymous):

Thanks Can you help me with another

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

ok Choose the graph below that represents the following system of inequalities: y greater than or equal to -3x + 1 y greater than or equal to 1 over 2x + 3

OpenStudy (texaschic101):

y >= -3x + 1 y int = (0,1) x int = (1/3,0) slope = -3 shaded above the line...solid line y >= 1/2x + 3 y int = (0,3) x int = (6,0) slope = 1/2 shaded above the line....solid line I think it is the 2nd one but let me check by taking 2 points in the shaded region (2,1) y >= -3x + 1 1 >= -3(2) + 1 1 >= -6 + 1 1 >= -5 (correct) Its the 2nd one

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

wait.....I think I messed up....I am going with the 1st one

OpenStudy (texaschic101):

there shaded above the line...so its the first one